OpenStudy (anonymous):

Let f(x) = x² + 6 and g(x)=x+8/x find (g*f)(-7) a)-55/7 b)384/7 c)295/49 d)63/55

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well f*g(x) will be \[\frac{(x^2 + 6)(x + 8)}{x}\] just substitute x = -7 and evaluate

OpenStudy (anonymous):

i got D as my answer is that right?

OpenStudy (campbell_st):

I got A (49 + 6)(-7 + 8)/-7 = (55 x 1)/-7 = - 55/7